You are here

Beam Delivery News

Primary tabs

TItle Date Excerpt Attached images
Beam delivery update 16 September 2014

The cyclotron is running and delivering ~105 µA to BL1A and ~95 µA to BL2C4. BL2A is just coming on following today’s shielding block move and stable-beam tuning to the ISAC yield station.

Beam delivery update: very little maintenance, if any, tomorrow [UPDATED] 15 September 2014

The cyclotron is running normally and delivering ~105 µA to BL1A, 9.8 µA to BL2A, and 90–95 µA to BL2C4. We expect to run like this until 06:00 tomorrow. Cyclotron operation may be interrupted briefly tomorrow morning for maintenance but at this point there is nothing planned. 

Beam delivery update: Weekend edition 12 September 2014

The cyclotron is running normally and delivering ~105 µA to BL1A, 9.8 µA to BL2A, and nA as needed to BL2C1 for PIF. PIF is scheduled to wrap up this evening, at which point STF operation on BL2C4 will resume; otherwise, we expect to run like this through the weekend.

Beam delivery update 11 September 2014

The cyclotron’s delivering ~105 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue overnight. 

Beam delivery update 10 September 2014

The cyclotron is currently running with ~20 µA in BL1A; the plan for the evening is to bring BL1A up to full current, then to bring BL2C4 back online. BL2A will remain off until tomorrow morning.

Beam delivery update: Recovery and the next ~48 hours 09 September 2014

After last night’s spontaneous start of a cryopanel defrost, the Cryogenics group came in to defrost and restart the tank cryogenics system overnight. It’s still not clear when beam might be injected – it could be anywhere between this evening and tomorrow morning. Operation over the next couple of days is still somewhat uncertain.

Beam delivery update [UPDATED] 08 September 2014

The cyclotron is still delivering nA as needed to BL1B, 9.8 µA to BL2A, and ~95 µA to BL2C4 for isotope production. PIF operation is scheduled to end tonight, but BL2A and 2C4 will continue to operate until the start of maintenance at 06:00 Wednesday.

Beam delivery update: Weekend edition [UPDATED] 05 September 2014

The cyclotron is delivering 9.8 µA to BL2A and ~95 µA to BL2C4, while BL1B has finished operation for the evening. PIF will resume tomorrow morning. Otherwise, we’ll continue like this through the weekend.

Beam delivery update 04 September 2014

The cyclotron is running much as it was last night, delivering nA to BL1B, 9.8 µA to BL2A, and ~95 µA to BL2C4. Operation is scheduled to carry on like this into the weekend.

Beam delivery update 03 September 2014

The cyclotron’s running normally, delivering nA as needed to BL1B for PIF, 9.8 µA to BL2A, and ~95 µA to BL2C4. We expect to run like this until 06:00 tomorrow morning, at which point BL2A will again be defined off to allow the experimenters to access the separator room to recover their target foil.

Pages

Subscribe to Group News